Physics Honors Section 84 Problem 8.18 Problem 8.18 While at the county fair, you decide to ride the Ferris wheel. Having eaten too many candy apples and elephant ears, you find the motion somewhat unpleasant. To take your mind off your stomach, you wonder about the motion of the ride. You estimate the radius of the big wheel to be 14 m and you use your watch to find that each loop around takes 23 s Part A What is yo Express y Submit Part B What is the Express yo Submit Part C What is the rate Express your
media%2F05f%2F05fbec68-2c45-4c74-840e-16
media%2F618%2F6183809c-ac1e-4a7a-8151-6b
Answer #1

Here,

r = 14 m

angular speed , w = 2pi/23 = 0.273 rad/s

part a)

speed = r * w

speed = 0.273 * 14 = 3.82 m/s

part B)

magnitude of acceleration = 3.82^2/14

magnitude of acceleration = 1.044 m/s^2

part C)

at the top of the ride

Wtop/FG = (9.8 - 1.044)/9.80

Wtop/FG = 0.893

part D)

at the bottom


Wbottom/FG = (9.8 + 1.044)/9.80

Wbottom/FG = 1.106